OpenStudy (anonymous):

-2x+3y=-3 -2x - 5y=-75

OpenStudy (anonymous):

Once again, multiply the second equation by -1 to get 2x+5y=75. Add the two equations up to get 8y=72. y=9. then substitute 9 into the first equation. -2x+3(9)=-3 -2x+27=-3 -2x=-30 x=15.

OpenStudy (anonymous):

put what you want to solve for

OpenStudy (anonymous):

for x and y

OpenStudy (anonymous):

I just did it for you.

OpenStudy (anonymous):

any questions?

OpenStudy (anonymous):

-2(15)+3y=-3 3y=27 y=9

OpenStudy (anonymous):

so y=72

OpenStudy (anonymous):

no. y=9.

OpenStudy (anonymous):

look at my first post, 2nd line.

OpenStudy (anonymous):

and x =15?

OpenStudy (anonymous):

yes it does.

OpenStudy (anonymous):

THANKSSSSSSSSSSSSSSSS!!!!!!!!!!! you were right

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):

=))))))))))))))

OpenStudy (anonymous):

can you help me with another?

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

-x - 4y=-37 -2x -5y=-53

OpenStudy (anonymous):

ok, so you would multiply the first equation by -2. and get 2x+8y=74. Do you get this?

OpenStudy (anonymous):

using that equation, add it to the other one to get 3y=21. Therefore, y=7. Substitute 7 for y in the first equation. -x-4(7)=-37. -x-28=-37 -x=9 x=-9.